Topics In Contemporary Math
Exponential Growth
5)The population of bison was almost hunted to extinction when the United States
expanded westward. The population decreased by 40% per year when the expansion
began in 1803. If the herds of bison originally contained 500,000 bison, how many were
left in 1820?

Answer by math_tutor2020(3816)


x = number of years
y = population in thousands

The gap from 1803 to 1820 is 1820-1803 = 17 years.


Plug in x = 17 to find the y value
y = a*b^x
y = 500*(1-0.40)^x
y = 500*(0.60)^x
y = 500*(0.60)^17
y = 0.084633 approximately

y is the population in thousands, so multiply that approximate result by 1000
1000*0.084633 = 84.633
That rounds to 85

Answer: 85